c=5/9(f-32), for f
Answer by nyc_function(2741)

Given c = 5/9(f-32), solve for f.

c=5/9(f-32) can also be written as c = 5(f - 32)/9.
Multiply each side by 9.
That gives us 9c = 5(f - 32).
Apply distributive rule on the right side.
9c = 5f - 160
Add 160 to the left side.
9c + 160 = 5f
Lastly, divide both sides 5.
(9c + 160)/5 = f